# Science Time Series: Deep Learning in Hydrology

### v1

@@ VII Conclusion and Future Work
Often, one thinks about deep learning models discovering hidden variables that control observables that one wishes to predict. In this paper, we explored CAMELS and Caravan, which together offer large datasets covering many (about 8000) spatial locations and a long time period (around 30 years). Further, there are multiple (6 in CAMELS and 39 in Caravan) observed dynamic streams and up to 209 static (exogenous) parameters for each location. Some of these exogenous and observed data would be natural inputs into a physics model for the catchments. A priori, it is not obvious which exogenous or even endogenous data should be used as they could implied by the large datasets used in the training and implicitly learned by the training. However, most likely, it is a mixed situation where the combination of observed and exogenous data will give the best results, with the exogenous data declining in importance as the observed dataset grows in size. We clearly show in this paper with the CAMELS and Caravan data that using exogenous data makes significant improvements in the model fits where both static physical quantities and dynamical mathematical functions have value as exogenous information. In future papers, we hope to understand the trade-off better as it varies in the nature and size of exogenous data. We also will present further study of different deep learning models using the compilation of hundreds of papers and around 100 models in [57, 58, 59, 60, 61]. We aim to study a range of time series problems to build a taxonomy so that benchmark sets can be built covering the essentially different cases [62]. Another interesting feature of science time series is that they naturally vary in magnitude by large factors, and this seems not to be very consistent with neural network activation functions at fixed values independent of the stream. We will present a study of new activation functions in the LSTM using a PRELU-like (Parametric Rectified Linear Unit, or PReLU) structure, which naturally deals with magnitude changes in stream values.
Acknowledgment
We would like to thank Gregor von Laszewski, Niranda Perera, and Alireza Jafari for their contributions and guidance. We gratefully acknowledge the partial support of DE-SC0023452: FAIR Surrogate Benchmarks Supporting AI and Simulation Research and the Biocomplexity Institute at the University of Virginia.

### v2

@@ 9 Conclusion and Future Work
Often, one thinks about deep learning models discovering hidden variables that control observables that one wishes to predict. In this paper, we demonstrate how including environmental factors and exogenous features can improve time series models’ ability to learn hidden variables. We explored CAMELS and Caravan, which together offer large datasets covering many (about 8000) spatial locations and a long time period (around 30 years). Further, there are multiple (6 in CAMELS and 39 in Caravan) observed dynamic streams and up to 209 static (exogenous) parameters for each location. Some of these exogenous and observed data would be natural inputs into a physics model for the catchments. A priori, it is not obvious which exogenous or even endogenous data should be used as they could be implied by the large datasets used in the training and implicitly learned by the training. However, most likely, it is a mixed situation where the combination of observed and exogenous data will give the best results, with the exogenous data declining in importance as the observed dataset grows in size. We clearly show in this paper with the CAMELS and Caravan data that using exogenous data makes significant improvements in the model fits where both static physical quantities and dynamical mathematical functions have value as exogenous information. We emphasize the importance of considering all available data in time series analyses. In future papers, we hope to understand the trade-off better as it varies in the nature and size of exogenous data. We also will present in-depth study of time series pattern and foundation models under varying hyperparameters, using the compilation of hundreds of papers and around 100 models in sci (2023); nix (n.d.); ddz (n.d.); Jin et al. (2023); Liang et al. (2024). We aim to study a range of time series problems to build a taxonomy so that benchmark sets can be built covering the essentially different cases Huang et al. (2019). Through a comprehensive model review, we will also examine univariate versus multivariate time series analysis. Another interesting feature of science time series is that they naturally vary in magnitude by large factors, and this seems not to be very consistent with neural network activation functions at fixed values independent of the stream. We will present a study of new activation functions in the LSTM using a PRELU-like (Parametric Rectified Linear Unit, or PReLU) structure, which naturally deals with magnitude changes in stream values.
Our LSTM model, data preprocessing, and model comparison code is available at https://github.com/JunyangHe/Hydrology. The processed data is available at He (2024).
{acks}
We would like to thank Gregor von Laszewski and Niranda Perera for their contributions and guidance.
{dci}
The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
{funding}
The author(s) disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work was partially supported by DE-SC0023452: FAIR Surrogate Benchmarks Supporting AI and Simulation Research and the Biocomplexity Institute at the University of Virginia.
{biogs}
Geoffrey Fox received a Ph.D. in Theoretical Physics from Cambridge University, where he was Senior Wrangler. He is now a Professor in the Biocomplexity Institute and Computer Science Department at the University of Virginia. He has an hindex of 90 with over 45,000 citations. He received the High-Performance Parallel and Distributed Computing (HPDC) Achievement Award and the ACM - IEEE CS Ken Kennedy Award for Foundational contributions to parallel computing in 2019. He is a Fellow of APS (Physics) and ACM (Computing) and works on the interdisciplinary interface between computing and applications.
Junyang He is an undergraduate student in computer science at the University of Virginia School of Engineering and Applied Sciences. He is an undergraduate research assistant under the supervision of Professor Geoffrey Fox. His research centers on time series analysis and large-scale spatiotemporal data modeling for scientific applications.
Alireza Jafari is a Ph.D. student in Computer Science at the University of Virginia and a Graduate Research Assistant under the supervision of Professor Geoffrey Fox. He earned his master’s degree in computer science, focusing on graph machine learning from the University of Tehran. Alireza’s research interests include machine learning, time series prediction, foundation models, graph neural networks, and AI applications in science.
Anushka Idamekorala is a Ph.D. student in Computer Science at the University of Virginia and a Graduate Research Assistant under the supervision of Professor Geoffrey Fox. He earned his bachelor’s degree in electronics and telecommunication engineering from the University of Moratuwa, Sri Lanka. Anushka’s research interests include machine learning, time series prediction, foundation models, and AI applications in science.
Ying-Jung Chen received the B.A. in geography from National Taiwan Normal University Taipei city, Taiwan, in 2008 and the M.A, Ph.D. degree in geography from the University of California, Santa Barbara Santa Barbara, CA, in 2018. She is currently pursuing her second master degree in computer science from Georgia Institute of Technology, Atlanta, GA. From 2018 to 2019, She was a joint post-doc Research Associate with the Applied Physics Laboratory and the e-Science Institute from the University of Washington, Seattle, Seattle, WA. Since 2020, She has been a Data and Applied Scientist/Machine Learning Engineer among different industrial sectors, including agriculture, finance,non-profit, and government. Her research interests include time-series analysis/forecasting, AI ethics and safety, remote sensing and geospatial applications. She has served on the program committee forClimate Change AI workshops at NeurIPS 2022-2024, ICLR 2023-2025, and datasafe workshops at AAAI 2025. Finally. she is the reviewer of the journal Water resources management, and Weather, climate, and society.